# Multifidelity linear regression for scientific machine learning from scarce data

### v1

@@ Abstract
Machine learning (ML) methods, which fit to data the parameters of a given parameterized model class, have garnered significant interest as potential methods for learning surrogate models for complex engineering systems for which traditional simulation is expensive. However, in many scientific and engineering settings, generating high-fidelity data on which to train ML models is expensive, and the available budget for generating training data is limited. ML models trained on the resulting scarce high-fidelity data have high variance and are sensitive to vagaries of the training data set. We propose a new multifidelity training approach for scientific machine learning that exploits the scientific context where data of varying fidelities and costs are available; for example high-fidelity data may be generated by an expensive fully resolved physics simulation whereas lower-fidelity data may arise from a cheaper model based on simplifying assumptions. We use the multifidelity data to define new multifidelity Monte Carlo estimators for the unknown parameters of linear regression models, and provide theoretical analyses that guarantee the approach’s accuracy and improved robustness to small training budgets. Numerical results verify the theoretical analysis and demonstrate that multifidelity learned models trained on scarce high-fidelity data and additional low-fidelity data achieve order-of-magnitude lower model variance than standard models trained on only high-fidelity data of comparable cost. This illustrates that in the scarce data regime, our multifidelity training strategy yields models with lower expected error than standard training approaches.
@@ 1 Introduction
Scientific and engineering decision-making rely on many-query analyses, in which a predictive model must be evaluated many times at different inputs, initializations, or parameters. Examples include optimization, control, and uncertainty quantification. Traditional high-fidelity models of complex scientific and engineering systems are often prohibitively expensive for this many-query setting, necessitating the development and use of computationally efficient surrogate models. Machine learning (ML) methods, which fit to data the parameters of a given parametrized model class, can learn extremely complex functional relationships from data [28, 37]. A growing body of literature uses such methods to learn surrogate models for engineering systems, for example using linear and kernel regressions [51, 55, 54, 44, 43, 65, 66, 46], as well as nonlinear regressions, for example using neural networks [63, 37, 26, 3, 35, 48, 39]. However, these works often take for granted the existence of (or ability to generate) a sufficiently large volume of high-fidelity training data needed to learn an accurate and robust model. This is a barrier to more widespread adoption of ML surrogate modeling methods in engineering and science because realistic budget constraints in these settings mean that only a limited amount of high-fidelity training data may be affordable. ML surrogate models trained on scarce data are highly sensitive to quirks of the data set and lead to less accurate and robust predictions [16], limiting trust in the learned models for use in high-consequence engineering and scientific application.
In this paper, we address the challenge of learning robust, accurate, and trusted ML surrogate models for engineering and scientific systems from scarce high-fidelity data by proposing a new multifidelity training approach to linear regression for scientific machine learning. Our approach exploits the scientific and engineering context in which data are often available from multiple sources/models with different levels of fidelity. For example, while high-fidelity data may come from a high-resolution multi-physics simulation of a three-dimensional system, lower-fidelity data may be available from simulations that use lower resolutions, make simplifying physics assumptions, or represent 1D or 2D approximations. We consider the training of ML models that are linear in the unknown model parameters that are fitted to data, i.e., the linear regression problem, and propose new multifidelity Monte Carlo estimators for the ML model parameters. The estimators combine a larger volume of cheaper-to-obtain low-fidelity data with a limited volume of expensive high-fidelity data to improve the robustness of the training process while remaining unbiased with respect to the high-fidelity distribution: this leads to multifidelity learned models that have lower expected error than models trained with only high-fidelity data of comparable cost. We will provide theoretical and numerical results demonstrating the efficacy of the multifidelity approach.
There are several bodies of work in multifidelity machine learning that combine models and/or data of varying fidelity. Some existing approaches use high- and low-fidelity data in separate training phases. One way to do this is to learn one model from high-fidelity data, and other models from low-fidelity data, and then to combine these separate learned models to issue multifidelity predictions [21, 24]. Another approach is multifidelity transfer learning, where low-fidelity data are first used to pre-train a model, and then high-fidelity data are subsequently used to refine the learned model parameters [57, 13, 31, 14, 62]. Other multifidelity learning approaches use all fidelities of data at once during training. Several works in this direction use high- and low-fidelity data to learn a correction or discrepancy for a provided low-fidelity model, and then combine the low-fidelity model with the learned correction model to issue multifidelity predictions [15, 25, 8, 33, 17, 1, 45, 67, 19]. These approaches can be viewed as hybrid approaches that mix a machine-learned correction with a low-fidelity model determined from first principles (e.g., linearization or simplified physics). Other approaches learn a low-fidelity model from low-fidelity data, and then learn a correction to this learned model: in some cases, the low-fidelity model is trained first and then the correction is trained [59, 36, 38], in other cases, the low-fidelity model is joined to the correction model and the training for both proceeds simultaneously in an ‘end-to-end’ training process [29, 1, 42].
In contrast to the aforementioned works, we propose a multifidelity training approach for linear regression models that uses all fidelities of data simultaneously to learn a single predictive model that is not a correction to an existing low-fidelity model. Our approach is simpler than approaches which require multiple training phases or linking multiple models in one end-to-end training phase. Other approaches which use all data simultaneously to learn a single model include multifidelity Gaussian process regression and data augmentation strategies. Multifidelity Gaussian process regression [4, 53, 40, 11] and co-kriging approaches [32, 34, 49, 52] fit parameters of Gaussian process regression models by exploiting the structure of the correlation between the data of varying fidelities in a Bayesian parameter update, where the choice of prior distribution for the learned parameters can have significant influence on the accuracy of the learned model predictions. In contrast, our approach is non-Bayesian, and thus does not require the imposition of prior information. Data augmentation strategies combine high- and low-fidelity data sets, potentially with different weights [59], or by mapping the high- and low-fidelity data to a shared feature space [59, 60] in which the model is then learned. Our approach is based on new multifidelity Monte Carlo estimators for the learned model parameters that go beyond data augmentation.
Multifidelity Monte Carlo (MFMC) and multilevel Monte Carlo (MLMC) are closely related methods that improve the robustness of Monte Carlo estimators by using samples from models of different fidelities and costs to reduce estimator variance. The two approaches are both based on control variate methods: a control variate is a random variable whose mean is (approximately) known, and which is correlated with the random variable of interest. Control variate methods exploit this correlation by adding a correction term to Monte Carlo estimators for statistics of the random variable of interest. In MFMC and MLMC, low-fidelity models serve as control variates. In MLMC, the low-fidelity models are assumed to belong to a model hierarchy with known convergence rates with respect to the underlying truth (e.g., via grid convergence of discretized PDEs), allowing convergence of the MLMC estimator to be proved [64, 23, 22]. In MFMC methods, more general low-fidelity models (without convergence rate assumptions) are considered [47, 50]. The MFMC mean estimator is unbiased, and has a lower standard error than a high-fidelity estimator of comparable cost if the low-fidelity models are correlated with the high-fidelity model and sufficiently cheap to evaluate [50]. In MLMC, samples are distributed among the models in order to achieve a desired error with respect to the underlying truth, whereas in MFMC the works [50, 56] propose optimal strategies for allocating a limited computational budget among different models to minimize the variance of the MFMC mean and variance estimators, respectively. MLMC ideas have been used for multifidelity machine learning in [21], which learns a hierarchy of neural networks of different levels and then combines these neural networks in an MLMC control variate framework to issue predictions, as well as in [7], which uses multilevel sequential Monte Carlo sampling to compute the parameters of a deep neural network in a Bayesian way. Another use of MLMC ideas in machine learning is in the stochastic estimation of gradients during the training process for neural networks, where multilevel estimators reduce the variance of the gradient estimator and can lead to improved convergence of the optimization [61, 20]. In contrast to these approaches, we propose new MFMC estimators for the unknown model parameters themselves in the linear regression setting.
Our contributions are the following:
@@
    We propose a new multifidelity machine learning approach for linear regression problems that can learn surrogate models for scientific and engineering systems from data of varying fidelities and costs.
@@
    Our approach proposes new multifidelity Monte Carlo estimators for the correlation between model inputs and output, and exploits the structure of linear regression problems in the scientific setting to propose multifidelity Monte Carlo estimates for the model parameters themselves.
@@
    We prove accuracy guarantees that show that our multifidelity estimators for the model parameters are unbiased with respect to high-fidelity estimators, and that the resulting model predictions are also unbiased.
  4. 4.
    We provide analysis of the variance of the learned model predictions that inform optimal choices of the coefficients of the control variates, which are hyperparameters of our method. These optimal choices improve the robustness of the multifidelity learned model to realizations of the data.
  5. 5.
    We provide numerical results on both an analytical example as well as a convection-diffusion-reaction model problem demonstrating that the proposed multifidelity approach learns from scarce data models of similar expected quality to standard learned models trained on orders-of-magnitude more high-fidelity data.
@@ 2.1 Problem formulation
by selecting $\beta\in\mathbb{R}^{d}$ such that $\hat{f}(Z;\beta)\approx Y^{(1)}$.
Ideally, we would like to find the parameter $\beta$ that minimizes the expected square error over the entire probability distribution, as follows:
@@
In practice, computing exact expectations with respect to the measure $\pi$ is often not possible, so the expectations in (3) are estimated from data. The standard approach uses a training data set $\{(x_{i},y_{i}^{(1)})\}_{i=1}^{n}$, where $y_{i}^{(1)}=f^{(1)}(z_{i})$ and $x_{i}=x(z_{i})$ for $z_{i}$ that are drawn i.i.d. from $\pi$. Denote by $X_{n}\in\mathbb{R}^{d\times n}$ the matrix whose $n$ columns are the feature data $x_{1},\ldots,x_{n}$, and denote by $Y_{n}^{(1)}\in\mathbb{R}^{n}$ the vector whose elements are the output data $y_{1}^{(1)},\ldots,y_{n}^{(1)}$. Then, a standard approximation to (3) is given by
@@
where $\hat{C}_{XX}^{(n)}=\frac{1}{n}X_{n}X_{n}^{\top}$ and $\hat{C}_{XY}^{(n)}=\frac{1}{n}X_{n}Y_{n}^{(1)}$ are $n$-sample estimates of the exact expected products $C_{XX}\equiv\mathbb{E}[XX^{\top}]$ and $C_{XY}\equiv\mathbb{E}[XY^{(1)}]$, respectively.
@@
where we assume we have the ability to compute $C_{XX}=\mathbb{E}[XX^{\top}]$ either exactly or with very many ($N\gg n$) input samples, but have only limited (small $n$) output samples with which to compute $\hat{C}_{XY}^{(n)}=\frac{1}{n}X_{n}Y_{n}^{(1)}$. Our analysis in Section 3.3 will assume $C_{XX}$ in (5) is known exactly.
Randomness of regression solution and non-robustness to scarce data. The estimated regression coefficients (5) are themselves random variables whose realizations depend on the training data. In scientific settings where computational resources are limited, and evaluating $f^{(1)}$ to obtain training data is expensive, we are often forced to work with small sample sizes $n$, which can lead to the estimated regression coefficients having high variance. Consider, for a fixed $z$, the conditional variance of the model:
@@
Because $\hat{\beta}^{(n)}$ estimated from scarce data (small $n$) has high variance, estimates from scarce data are likely to be far from the $\beta^{*}$ given by (2) which minimizes the expected square error of the model. From eq. 6, we see that this in turn leads to learned model predictions that are less accurate in expectation because they are likely to be far from predictions of the ideal model defined by $\beta^{*}$. The goal of this paper is to introduce a multifidelity training strategy that uses data from both cheap low-fidelity models and the expensive high-fidelity model to reduce the variance of the estimated regression coefficients $\hat{\beta}$ while guaranteeing unbiasedness of the estimators. This in turn will lead to learned model predictions with lower variance and thus higher expected accuracy.
2.2 Multifidelity Monte Carlo
We now summarize the multifidelity Monte Carlo (MFMC) approach to estimating expected values, which is based on the control variate variance reduction technique for Monte Carlo methods. Recall that the standard Monte Carlo estimator of $\mathbb{E}[f^{(1)}(Z)]$ is given by
@@ 2.2 Multifidelity Monte Carlo
where $\{z_{i}\}_{i=1}^{n}$ are i.i.d. realizations of the input random variable $Z$. This mean estimator $\hat{\mu}_{n}$ is unbiased: $\mathbb{E}[\hat{\mu}_{n}]=\mathbb{E}[f^{(1)}(Z)]$, and its variance is $\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}]=\frac{1}{n}\operatorname{\mathbb{V}ar}[f^{(1)}(Z)]$.
A control variate is another random variable $W$ with mean $\mathbb{E}[W]=\mu_{W}$, which defines the following control variate estimator:
$\hat{\mu}_{n}^{\rm CV}=\frac{1}{n}\sum_{i=1}^{n}f^{(1)}(z_{i})+\alpha\left(\mu_{W}-\frac{1}{n}\sum_{i=1}^{n}w_{i}\right),$ (7)
where $\alpha\in\mathbb{R}$ is called the control variate coefficient. If $\mu_{W}$ is known exactly, the formula (7) defines an exact control variate. If we replace $\mu_{W}$ in (7) with a sample estimate, then the resulting estimator is an approximate control variate. Both the exact and approximate estimators (7) are unbiased, $\mathbb{E}[\hat{\mu}_{n}^{\rm CV}]=\mathbb{E}[f^{(1)}(Z)]$. The variance of the exact control variate estimator is given by
$\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}^{\rm CV}]=\frac{\operatorname{\mathbb{V}ar}[f^{(1)}(Z)]}{n}+\alpha^{2}\frac{\operatorname{\mathbb{V}ar}[W]}{n}-2\alpha\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),W]}{n}.$
@@
$\alpha^{*}=\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),W]}{\operatorname{\mathbb{V}ar}[W]}.$
@@
$\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}^{\rm CV}]=\frac{1}{n}\left(\operatorname{\mathbb{V}ar}[f^{(1)}(Z)]-\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),W]^{2}}{\operatorname{\mathbb{V}ar}[W]}\right).$
Thus, the exact control variate estimator has lower variance than the standard Monte Carlo estimator if the control variate $W$ is correlated with $f^{(1)}(Z)$: the better the correlation, the greater the variance reduction. Approximate control variate estimators have slightly different variance expressions and associated variance reduction conditions, but the main idea that improved correlation leads to greater variance reduction remains the same, see [50] for details.
The main idea of multifidelity Monte Carlo (MFMC) is to use low-fidelity models to define approximate control variates, i.e., let $f^{(2)}:\mathbb{R}^{d}\to\mathbb{R}$ be a lower-fidelity model for the same system described by $f^{(1)}$, and let the control variate be given by $W=f^{(2)}(Z)$. We now present the MFMC method introduced and analyzed in [50], which generalizes to a hierarchy of $K\geq 2$ models $f^{(1)},\ldots,f^{(K)}:\mathbb{R}^{d}\to\mathbb{R}$. Let $w_{k}$ denote the cost of model $f^{(k)}$ and let
@@
$\hat{\mu}_{\rm MF}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}f^{(1)}(z_{i})+\sum_{k=2}^{K}\alpha_{k}\left(\frac{1}{m_{k}}\sum_{i=1}^{m_{k}}f^{(k)}(z_{i})-\frac{1}{m_{k-1}}\sum_{i=1}^{m_{k-1}}f^{(k)}(z_{i})\right),$ (9)
where $m_{1}<m_{2}<\cdots<m_{K}$ are the number of model evaluations for each of the $K$ models, and $\{z_{i}\}_{i=1}^{m_{K}}$ are i.i.d. draws of the input random variable as before. One way to interpret (9) is to view the control variate based on lower-fidelity model $k$ as a higher-sample correction to the estimator based on fewer samples of the higher-fidelity models $1,\ldots,(k-1)$.
The MFMC estimator is unbiased, i.e., $\mathbb{E}[\hat{\mu}_{\rm MF}]=\mathbb{E}[f^{(1)}(Z)]$, and the MFMC estimator has a lower variance than the standard Monte Carlo estimator of the same cost, i.e., if $\sum_{k=1}^{K}m_{k}w_{k}=nw_{1}$, then
$\operatorname{\mathbb{V}ar}[\hat{\mu}_{\rm MF}]<\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}],$
provided that the models satisfy certain conditions on their relative costs and correlations (see [50, Corollary 3.5]). The work [50] provides a model selection algorithm that selects models so that these conditions are satisfied. Additionally, the work [50] provides an optimal allocation of a limited computational budget among the $K$ different models that minimizes $\operatorname{\mathbb{V}ar}[\hat{\mu}_{\rm MF}]$: that is, the optimal assignment of $m_{1},\ldots,m_{K}$ such that $\sum_{k=1}^{K}m_{k}w_{k}=p$, where $p$ is the computational budget. This optimal assignment is given by

### v2

@@ Abstract
Machine learning (ML) methods, which fit to data the parameters of a given parameterized model class, have garnered significant interest as potential methods for learning surrogate models for complex engineering systems for which traditional simulation is expensive. However, in many scientific and engineering settings, generating high-fidelity data on which to train ML models is expensive, and the available budget for generating training data is limited, so that high-fidelity training data are scarce. ML models trained on scarce data have high variance, resulting in poor expected generalization performance. We propose a new multifidelity training approach for scientific machine learning via linear regression that exploits the scientific context where data of varying fidelities and costs are available: for example, high-fidelity data may be generated by an expensive fully resolved physics simulation whereas lower-fidelity data may arise from a cheaper model based on simplifying assumptions. We use the multifidelity data within an approximate control variate framework to define new multifidelity Monte Carlo estimators for linear regression models. We provide bias and variance analysis of our new estimators that guarantee the approach’s accuracy and improved robustness to scarce high-fidelity data. Numerical results demonstrate that our multifidelity training approach achieves similar accuracy to the standard high-fidelity only approach with orders-of-magnitude reduced high-fidelity data requirements.
@@ 1 Introduction
Scientific and engineering decision-making rely on many-query analyses, in which a predictive model must be evaluated many times at different inputs, initializations, or parameters. Examples include optimization, control, and uncertainty quantification. Traditional high-fidelity models of complex scientific and engineering systems are often prohibitively expensive for this many-query setting, necessitating the development and use of computationally efficient surrogate models. Machine learning (ML) methods, which fit to data the parameters of a given parametrized model class, can learn extremely complex functional relationships from data [35, 44]. A growing body of literature uses such methods to learn surrogate models for engineering systems, for example using linear and kernel regressions [56, 61, 60, 50, 51, 73, 74, 53], as well as nonlinear regressions, for example using neural networks [71, 44, 33, 3, 42, 54, 46]. These works often take for granted the existence of (or ability to generate) a sufficiently large volume of high-fidelity training data for learning an accurate and robust model. This is a barrier to more widespread adoption of ML surrogate modeling methods in engineering and science because realistic budget constraints in these settings mean that high-fidelity data are usually scarce, due to the expense of obtaining data through expensive simulations or experiments. ML surrogate models trained on scarce data are sensitive to quirks of the data set and lead to less accurate and robust predictions [18], limiting trust in the learned models for use in high-consequence engineering and scientific application.
One way to address the data scarcity challenge is through multifidelity scientific machine learning, which seeks to learn models from a combination of scarce high-fidelity data and more abundant lower-cost, lower-fidelity data, which are often available in scientific and engineering contexts. For example, while high-fidelity data may come from a high-resolution multi-physics simulation of a three-dimensional system, lower-fidelity data may be available from simulations that use lower resolutions or make simplifying physics assumptions. We first review approaches in which several separate (learned) models are combined to produce a single multifidelity model. These approaches include works that learn separate models from high- and low-fidelity data and them combine these models to issue multifidelity predictions [24, 30], as well as hybrid approaches which use high- and low-fidelity data to learn a correction or discrepancy for a provided physics-based low-fidelity model, and then combine the low-fidelity model with the learned correction model to issue multifidelity predictions [17, 31, 9, 40, 19, 1, 52, 75, 22]. One can also learn a low-fidelity model from low-fidelity data, and then learn a correction term: in some cases, the low-fidelity model is trained first and then the correction term is trained [43, 45, 67]; in other cases, the low-fidelity model is joined to the correction model and the training for both proceeds simultaneously in an ‘end-to-end’ training process [36, 1, 49]. These approaches which jointly train low-fidelity and correction models can also be viewed as belonging to another family of multifidelity machine learning methods in which data of varying fidelity are combined during the training process to learn the parameters of a single multifidelity model. These approaches include transfer learning methods, which first use lower-fidelity data to pre-train a model, and then use high-fidelity data to refine estimates of a subset of the model parameters [63, 15, 38, 16, 70], as well as data augmentation strategies in which high- and low-fidelity data sets are combined, potentially with different weights [67], or by mapping the high- and low-fidelity data to a shared feature space [67, 68] in which the model is then learned. Bayesian learning methods provide another way to combine high- and low-fidelity data to train a single model: in multifidelity Gaussian process regression [5, 59, 47, 12] and co-kriging approaches [39, 41, 55, 58], the parameters of Gaussian process regression models are fit to multifidelity data by exploiting the structure of the correlation between the data of varying fidelities in a Bayesian parameter update. The works [29, 28] formulate a more general Bayesian multifidelity learning approach that can accommodate non-Gaussian distributions.
In this work, we propose a new multifidelity scientific machine learning approach for linear regression. We note that linear regression models, while comparatively simpler than nonlinear regression models such as neural networks, are a large class of models that can include models that are linear in features that are arbitrarily nonlinear in the model inputs, including features that are defined by neural networks [4, 53]. Our approach is based on multifidelity control variate methods for variance reduction of Monte Carlo estimators, which combine high-fidelity estimators computed using a small sample size with a correction term based on a larger set of low-fidelity samples. This correction term is formulated in a way that preserves unbiasedness of the control variate estimator with respect to the high-fidelity statistics, in contrast to some of the aforementioned correction approaches which do introduce bias. Multifidelity control variate methods include Multilevel Monte Carlo (MLMC) [72, 25, 26], Multifidelity Monte Carlo (MFMC) [56, 62], generalized approximate control variates (ACV) [27], and Multilevel Best Linear Unbiased Estimators (MLBLUE) [65, 66, 13, 20]. These approaches differ in the types of low-fidelity models they consider, and in the structure of their low-fidelity correction term(s). In contrast to multifidelity Bayesian inference methods like co-kriging, multifidelity control variate estimators are frequentist methods, although connections between the two approaches have been shown [29]. Multifidelity control variate ideas have been previously used for machine learning in several different ways: the work [24] learns a hierarchy of neural networks of different levels and then combines these neural networks in an MLMC control variate framework to issue predictions. The work [8] uses multilevel sequential Monte Carlo sampling to compute the parameters of a deep neural network in a Bayesian way, while in [20], the authors extend MLBLUE to the estimation of covariance matrices in the context of variational data assimilation problems. In the previously mentioned Bayesian multifidelity learning approaches [29, 28], the multifidelity Bayesian posterior mean can be shown to be a control variate estimator. Another use of MLMC ideas in machine learning is in the control variate estimation of gradients in gradient-based optimization for training neural networks [69, 23] or optimization under uncertainty [17].
To the best of our knowledge, this work presents the first use of multifidelity control variates for surrogate model learning via linear regression. Our contributions are the following:
@@
    We propose new multifidelity control variate estimators for the unknown parameters of linear regression models. Our overall approach is generally applicable to the use of any multifidelity control variate estimator, including ACV and MLBLUE estimators, but we specifically propose new MFMC estimators for linear regression for clarity of exposition.
@@
    We provide bias and variance analysis of the proposed MFMC estimators for linear regression that (i) guarantees unbiasedness of the multifidelity estimators for the unknown model parameters as well as the resulting multifidelity learned model predictions and (ii) informs optimal choices of hyperparameters for the method which improve the accuracy and robustness of our approach.
@@
    We conduct numerical experiments on both an analytical example as well as a convection-diffusion-reaction model problem. Our numerical results demonstrate that models learned from scarce high-fidelity data using our multifidelity approach have similar accuracy to standard learned models trained on orders-of-magnitude more high-fidelity data.
@@ 2 Background
We introduce the linear regression problem we consider in Section 2.1, followed by background on multifidelity control variate estimators in Section 2.2.
@@ 2.1 Problem formulation
by selecting $\beta\in\mathbb{R}^{d}$ such that $\hat{f}(Z;\beta)\approx Y^{(1)}$. Ideally, we would like to find the parameter $\beta$ that minimizes the expected square error over the entire probability distribution, as follows:
@@
In practice, computing exact expectations with respect to the measure $\pi$ is often not possible, so the expectations in (3) are estimated from data. The standard approach uses a training data set $\{(x_{i},y_{i}^{(1)})\}_{i=1}^{n}$, where $y_{i}^{(1)}=f^{(1)}(z_{i})$ and $x_{i}=x(z_{i})$ for $z_{i}$ that are drawn i.i.d. from $\pi$. Denote by $X_{n}\in\mathbb{R}^{d\times n}$ the matrix whose $n$ columns are the feature data $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$, and denote by $Y_{n}^{(1)}\in\mathbb{R}^{n}$ the vector whose elements are the output data $y_{1}^{(1)},\ldots,y_{n}^{(1)}$. Then, a standard approximation to (3) is given by
@@
where $\hat{C}_{XX}^{(n)}=\frac{1}{n}X_{n}X_{n}^{\top}$ and $\hat{c}_{XY}^{(n)}=\frac{1}{n}X_{n}Y_{n}^{(1)}$ are $n$-sample estimates of the exact expected products $C_{XX}\equiv\mathbb{E}[XX^{\top}]$ and $C_{XY}\equiv\mathbb{E}[XY^{(1)}]$, respectively. However, although $\hat{C}_{XX}^{(n)}$ is an unbiased estimator for $C_{XX}$, $(\hat{C}_{XX}^{(n)})^{-1}$ inverse is only an asymptotically unbiased estimator for $C_{XX}^{-1}$ [32], so $\tilde{\beta}^{(n)}$ is a biased estimator for $\beta^{*}$, particularly in the scarce data setting when $n$ is small.
@@
where we assume we have the ability to compute $C_{XX}=\mathbb{E}[XX^{\top}]$ either exactly or with very many ($N\gg n$) input samples, but have only limited (small $n$) output samples with which to compute $\hat{c}_{XY}^{(n)}=\frac{1}{n}X_{n}Y_{n}^{(1)}$. The estimator in (5) is an unbiased estimator for $\beta^{*}$, even when $n$ is small, and our analysis in Section 3.3 will rely on this fact.
Randomness of regression solution and non-robustness to scarce data. The estimated regression coefficients in (5) are themselves random variables whose realizations depend on the training data. In scientific settings, where computational resources are limited and evaluating $f^{(1)}$ to obtain training data is expensive, we are often forced to work with small sample sizes $n$, which can lead to the estimated regression coefficients having high variance. Consider, for a fixed $z$, the conditional variance of the model:
@@
Since $\hat{\beta}^{(n)}$ estimated from scarce data (small $n$) has high variance, estimates from scarce data are likely to be far from the $\beta^{*}$ given by (2). From eq. 6, we see that this in turn leads to learned model predictions that are less accurate in expectation because they are likely to be far from predictions of the ideal model defined by $\beta^{*}$. The goal of this paper is to introduce a multifidelity training strategy that uses data from both cheap low-fidelity models and the expensive high-fidelity model to reduce the variance of the estimated regression coefficients $\hat{\beta}$ while guaranteeing unbiasedness of the estimators. This in turn will lead to learned model predictions with lower variance and thus higher expected accuracy.
2.2 Multifidelity Control Variate Estimators
We now introduce the multifidelity control variate approach to reducing estimator variance. Recall that the standard Monte Carlo estimator of $\mathbb{E}[f^{(1)}(Z)]$ is given by
@@ 2.2 Multifidelity Control Variate Estimators
where $\{z_{i}\}_{i=1}^{n}$ are i.i.d. realizations of the input random variable $Z$. This mean estimator $\hat{\mu}_{n}$ is unbiased, i.e., $\mathbb{E}[\hat{\mu}_{n}]=\mathbb{E}[f^{(1)}(Z)]$, and its variance is $\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}]=\frac{1}{n}\operatorname{\mathbb{V}ar}[f^{(1)}(Z)]$.
A control variate is another random variable $U$ with mean $\mathbb{E}[U]=\mu_{u}$, which defines the following control variate estimator:
$\hat{\mu}_{n}^{\rm CV}=\frac{1}{n}\sum_{i=1}^{n}f^{(1)}(z_{i})+\alpha\left(\mu_{u}-\frac{1}{n}\sum_{i=1}^{n}u_{i}\right),$ (7)
where $\alpha\in\mathbb{R}$ is called the control variate coefficient and $u$ is a realization of $U$. If $\mu_{u}$ is known exactly, the formula (7) defines an exact control variate. If we replace $\mu_{u}$ in (7) with a sample estimate, then the resulting estimator is an approximate control variate. Both the exact and approximate estimators (7) are unbiased, $\mathbb{E}[\hat{\mu}_{n}^{\rm CV}]=\mathbb{E}[f^{(1)}(Z)]$. The variance of the exact control variate estimator is given by
$\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}^{\rm CV}]=\frac{\operatorname{\mathbb{V}ar}[f^{(1)}(Z)]}{n}+\alpha^{2}\frac{\operatorname{\mathbb{V}ar}[U]}{n}-2\alpha\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),U]}{n}.$
@@
$\alpha^{*}=\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),U]}{\operatorname{\mathbb{V}ar}[U]}.$
@@
$\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}^{\rm CV}]=\frac{1}{n}\left(\operatorname{\mathbb{V}ar}[f^{(1)}(Z)]-\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),U]^{2}}{\operatorname{\mathbb{V}ar}[U]}\right).$
Thus, the exact control variate estimator has lower variance than the standard Monte Carlo estimator if the control variate $U$ is correlated with $f^{(1)}(Z)$: the better the correlation, the greater the variance reduction. Approximate control variate estimators have more complex variance expressions and associated variance reduction conditions, but the main idea that improved correlation leads to greater variance reduction remains the same, see e.g., [57, 27] for details.
The main idea of multifidelity control variate estimation is to use low-fidelity models to define approximate control variates, i.e., let $f^{(2)}:\mathbb{R}^{d}\to\mathbb{R}$ be a lower-fidelity model for the same system described by $f^{(1)}$, and let the control variate be given by $U=f^{(2)}(Z)$. We can generalize this idea to a hierarchy of $K\geq 2$ models $f^{(1)},\ldots,f^{(K)}:\mathbb{R}^{d}\to\mathbb{R}$ by adding additional approximate control variates based on each model to the estimator (7). All multifidelity control variate strategies are variations on this theme: for clarity of exposition, we now introduce only the multifidelity Monte Carlo (MFMC) [57] approach. MFMC is the focus of Section 3, which introduces our method, and Section 4, which provides numerics. We emphasize that other multifidelity control variate strategies, such as ACV [27] and MLBLUE [65, 13] estimators, can be substituted everywhere an MFMC estimator is used, and we discuss key differences between MFMC and ACV and MLBLUE where appropriate.
Recall that we consider $K\geq 2$ models $f^{(1)},\ldots,f^{(K)}:\mathbb{R}^{d}\to\mathbb{R}$. Let $w_{k}$ denote the cost of model $f^{(k)}$ and let
@@
$\hat{\mu}^{\rm MF}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}f^{(1)}(z_{i})+\sum_{k=2}^{K}\alpha_{k}\left(\frac{1}{m_{k}}\sum_{i=1}^{m_{k}}f^{(k)}(z_{i})-\frac{1}{m_{k-1}}\sum_{i=1}^{m_{k-1}}f^{(k)}(z_{i})\right),$ (9)
where $m_{1}<m_{2}<\cdots<m_{K}$ are the number of model evaluations for each of the $K$ models, and $\{z_{i}\}_{i=1}^{m_{K}}$ are i.i.d. draws of the input random variable as before. One way to interpret (9) is to view the control variate based on the $k$-th lower-fidelity model as a higher-sample correction to the estimator based on fewer samples of the 1st-through-$(k-1)$-th higher-fidelity models.
The MFMC estimator is unbiased, i.e., $\mathbb{E}[\hat{\mu}^{\rm MF}]=\mathbb{E}[f^{(1)}(Z)]$, and the MFMC estimator has a lower variance than the standard Monte Carlo estimator of the same cost, i.e., if $\sum_{k=1}^{K}m_{k}w_{k}=nw_{1}$, then
$\operatorname{\mathbb{V}ar}[\hat{\mu}^{\rm MF}]\leq\operatorname{\mathbb{V}ar}[\hat{\mu}_{n}],$
provided that the models satisfy certain conditions on their relative costs and correlations: loosely speaking, low-fidelity models must have sufficiently low cost to justify their low correlation (see [57, Corollary 3.5] for details). The work [57] provides a model selection algorithm that selects models so that these conditions are satisfied. Additionally, the work [57] shows that the optimal choice of the coefficients $\alpha_{k}$ is given by $\alpha_{k}=\frac{\rho_{1k}\sigma_{1}}{\sigma_{k}}$, and also provides an optimal allocation of a limited computational budget among the $K$ different models that minimizes $\operatorname{\mathbb{V}ar}[\hat{\mu}^{\rm MF}]$: that is, the optimal assignment of $m_{1},\ldots,m_{K}$ such that $\sum_{k=1}^{K}m_{k}w_{k}=p$, where $p$ is the computational budget. This optimal assignment has an analytical solution and is given by